OpenStudy (anonymous):

i need help determining the slope of the line that passes through the points at (-1,5)and (-4,10) pleasssse help its my last question and im stuck

OpenStudy (anonymous):

slope m = (10-5)/(-4-(-1))

OpenStudy (anonymous):

m=5/(-3)=-5/3

OpenStudy (seamus58):

slope=(5-10)/((-1)-(-4))=-5/3

OpenStudy (anonymous):

thanks so much -5/3 is right thank you thank you thank you